OpenStudy (anonymous):

In a class, except 18 all are above 50 years. 15 are below 50 years of age. How many people are there

OpenStudy (anonymous):

need help

Parth (parthkohli):

15 are below 50 years of age. Except 18 are all above 50 years. That means, that the 18 can be equal to 18. 18 - 15 = 3 3 people are 50 years old. Under 50 + Equal to 50 + more than 50 15 + 3 + 15 => 33

Parth (parthkohli):

I'm not sure, I might need a confirmation from the experienced users.

OpenStudy (lgbasallote):

how did you get the second 15?

Parth (parthkohli):

3 people are 50 years old in that 18 people. 18 - 3 = 15.. Get it?

OpenStudy (lgbasallote):

yeah..15 are below 50...3 are 50...how did you get that 15 are above 50 /:)

Parth (parthkohli):

Except 18, all are ABOVE 50 ----> More than 50. 15 are below 50. So, where are those 3 people? They must be equal to 50 years, right?

OpenStudy (anonymous):

i'd say 17

OpenStudy (lgbasallote):

lol your logic is so screwed up @ParthKohli =)) except 18 all are above 18...that means 18 people are less than or equal to 50..

OpenStudy (anonymous):

 It cant be 18 and you cant have 17 and a half a person or quarter person you must have 17

OpenStudy (lgbasallote):

since 15 are below....then 3 are equal to 50

Parth (parthkohli):

Yes, that's why I said..

Parth (parthkohli):

@lgbasallote Do you mean that my logic is screwed up or what?

OpenStudy (lgbasallote):

where did you get 15 are above 50 then

OpenStudy (anonymous):

or maybe 21

OpenStudy (lgbasallote):

actually this question makes no sense because there were no indications how many people are above 50...either it was written incorrectly or just not enough info is given

OpenStudy (anonymous):

no sir the question is given in the same format... From yesterday I have been stryggling to get the ans. @lgbasallote

OpenStudy (anonymous):

i agree with you ! there is no info for above 50

Parth (parthkohli):

lgba = sir @lgbasallote your thoughts?

OpenStudy (anonymous):

Let assume that the no. of people above 50 is k and the no of people is x => 18 + k = x so the no. of people is (18+k) where k is a positive integer

OpenStudy (anonymous):

thanks guys...

OpenStudy (anonymous):

"In a class, except 18 all are above 50 years." - You mean that 18 are below or equal to 50 yrs So, substracting, we get : 18 - 15 = 3 ... 3 are equal to 50 yrs However, we can't proceed any further that this ...The "all above 50" mentioned in the first line is maybe for nothing. So, i'll say the total is imply 18

OpenStudy (anonymous):

3 are above or equal 50 15 below 50 that was given there fore there are 21 peeps

OpenStudy (anonymous):

@timo86m , except 18, all are above 50.

OpenStudy (anonymous):

there are 18 not = to 50 therefore 18+3 that are equal to 50 or above is 21